How much of an alloy that is 20% copper should be mixed with 300 oz of an alloy that is 80% copper to get an alloy that is 40% copper?

Answer by robertb(5830)

0.20x + 0.80*300 = 0.40*(x + 300)
==> 0.20x + 240 = 0.40x + 120
==> 120 = 0.20x
==> x = 600 oz of the 20% copper alloy should be added.